# ros2_fanuc_interface: Design and Evaluation of a Fanuc CRX Hardware Interface in ROS2

### v1

@@ Abstract
This paper introduces the ROS2 control and the Hardware Interface (HW) integration for the Fanuc CRX- robot family. It explains basic implementation details and communication protocols, and its integration with the Moveit2 motion planning library. We conducted a series of experiments to evaluate relevant performances in the robotics field. We tested the developed ros2_fanuc_interface for four relevant robotics cases: step response, trajectory tracking, collision avoidance integrated with Moveit2, and dynamic velocity scaling, respectively. Results show that, despite a non-negligible delay between command and feedback, the robot can track the defined path with negligible errors (if it complies with joint velocity limits), ensuring collision avoidance. Full code is open source and available at https://github.com/paolofrance/ros2_fanuc_interface.
@@ I Introduction
The Robot Operating System (ROS) eco-system is the de-facto standard to control robots in the research field and is growing interest from the industrial world. ROS and ROS2 allow for seamless integration of state-of-the-art motion planning, vision, and control algorithms, acting as a middleware for modular software development and deployment. While initially widespread mainly in the academia, ROS2 has now reached a level of maturity that makes it usable in industrial applications [1].
@@ II Preliminaries
When developing a hardware interface from scratch, it is essential to understand how a remote computer communicates with the robot controller. The Fanuc CRX robots are manged by the Fanuc R30iB Mini Plus controller, which allows for remote commands to the robot through two modalities, Remote Motion Interface (RMI) and Ethernet/IP (EIP).
@@ III ROS2-Fanuc Hardware Interface
This section describes the main components of the ROS2 Fanuc interface; namely, how we use Fanuc EIP interface to communicate with the robot and how the robot executes motion commands received from the external computer.
@@ III-A Fanuc Ethernet/IP
In the proposed control architecture, a ROS2-based driver running on an external computer communicates with the robot by writing desired target poses into PR[1], at a frequency specified by the ros2_control hardware interface. The motion execution logic is implemented in a TP program on the Fanuc controller according to the value stored into PR[1].
@@ III-B TP program
The combination of TP program and the ROS2 hardware interface works as follows:
@@
    If R[1]==1, the TP program continuously commands the robot to move to the target joint position avaiable in PR[1]. Using a CNT=100% motion instruction enables smooth trajectory execution.
This register-based if-then-else structure of the TP program can be extended to use other registers to commands other non-motion operations. For example, one may use R[2] to command the opening/closing of a gripper through a digital input signal. Several services, such as the gripper ones in the fanuc_srvs, work according to this principle.
@@ IV Performance evaluations
To check the performance and usability of the developed hardware interface, we conducted experiments for four relevant tasks in industrial robotics: reaching a target point, tracking a trajectory, executing a collision-free trajectory, and dynamic velocity scaling. Experiments are performed with a Fanuc CRX-25iA communicating via Ethernet with a remote PC at a frequency of $25Hz$. Position and velocity measurements were recorded at the ROS2 controllers level, which means they include communication latency and ROS2 possible computational overheads. Comparable results were observed with a CRX-10iA_L and CRX-20iA_L robots.
@@ IV-B Trajectory tracking
To measure the tracking capabilities of the hardware interface, we streamed sinusoidal reference trajectories, $ref(t)=A\sin(2\pi ft)$, with amplitude $A=30\degree$ and frequencies $f\in\{0.1,0.25,0.5\}\,Hz$. Also for this experiment, we requested the robot to move only the first joint. We evaluate the tracking error and the path following error.
@@
Figure 3: Results of the step resposes (top), trajectory tracking (middle), and path following (bottom).
@@ IV-B2 Path following error
In this case, for the low-frequency cases ($f=\{0.1,0.25\}Hz$), the errors dramatically decrease, with averages below $1\degree$. As for the higher frequency, $f=0.5Hz$, the velocity of the joint saturates, and the resulting error remains large, leading to poor path following. The maximum error is measured when the command changes from zero to $ref(t)$. In this case, the robot should have a non-zero velocity and must adhere to physical constraints. The path errors are in in the bottom rows of Table II and in Figure 3g, 3h, and 3i.
@@ IV-C Obstacle avoidance
We assessed the correct execution of collision-free trajectories via with Moveit2 [10] through the experiment shown in Figure 1. The robot has to move from an initial to a goal configuration by avoiding a box-shape obstacle. The trajectory is computed with the BiTRRT motion planner, with a velocity scaling of 0.1. The executed trajectory is in Figure 4. The plots confirm the results obtained in the previous experiments in terms of trajectory tracking and path following.
@@ IV-D Dynamic velocity scaling
We used the proposed hardware interface to implement online trajectory scaling for safe human-robot collaboration and collaborative transportation of objects. These experiments used a custom ROS2 controller, named scaled_fjt_controller, based on [11, 12] and available here [13].
Detailed experiments for velocity scaling based can be found in [14], where the proposed ros2_fanuc_interface is used in combination with the scaled_fjt_controller to adapt trajectory execution, based on interaction force dynamically. An example is in Figure 5, which shows the motion of the first joint when it is dynamically adjusted in real time. After about one second, the speed override (dashed line) drops to a value close to zero. The command slope flattens for about one second and then starts growing again because of the high speed override. A similar behavior is visible in the interval between 6 and 8 seconds.
Figure 5: Command and feedback signals with dynamic velocity scaling during trajectory execution. The velocity is modulated by an external force sensor.
@@ IV-E Limitations
The ros2_fanuc_interface does not enforce joint velocity limits. This means that large velocity commands may be clamped by the Fanuc controller. For this reason, careful attention must be placed on the robot controller’s settings, as the Fanuc controller may enforce Cartesian speed limits of around 250 mm/s in the collaborative mode. This may lead to unexpected joint speed clamping during execution.

### v2

@@ Abstract
This paper introduces the ROS2 control and hardware interface integration for the Fanuc CRX robot family. It explains basic implementation details and communication protocols, and its integration with the Moveit2 motion planning library. We conducted a series of experiments to evaluate relevant performances in the robotics field. We tested the developed ros2_fanuc_interface for four relevant robotics cases: step response, trajectory tracking, collision avoidance integrated with Moveit2, and dynamic velocity scaling, respectively. Results show that, despite a non-negligible delay between command and feedback, the robot can track the defined path with negligible errors (if it complies with joint velocity limits), ensuring collision avoidance. Full code is open source and available at https://github.com/paolofrance/ros2_fanuc_interface.
@@
This work is currently under review in a IEEE journal
@@ I Introduction
The Robot Operating System (ROS) ecosystem is the de facto standard to control robots in the research field and is growing in interest from the industrial world. ROS and ROS2 allow for seamless integration of state-of-the-art motion planning, vision, and control algorithms, acting as a middleware for modular software development and deployment. While initially widespread mainly in academia, ROS2 has now reached a level of maturity that makes it usable in industrial applications [1].
@@ II Preliminaries
When developing a hardware interface from scratch, it is essential to understand how a remote computer communicates with the robot controller. The Fanuc CRX robots are managed by the Fanuc R30iB Mini Plus controller, which allows for remote commands to the robot through two modalities, Remote Motion Interface (RMI) and Ethernet/IP (EIP).
@@ III ROS2-Fanuc Hardware Interface
This section describes the main components of the ROS2 Fanuc interface; namely, how we use the Fanuc EIP interface to communicate with the robot and how the robot executes motion commands received from the external computer.
@@ III-A Fanuc Ethernet/IP
In the proposed control architecture, a ROS2-based driver running on an external computer communicates with the robot by writing desired target poses into PR[1], at a frequency specified by the ros2_control hardware interface. The motion execution logic is implemented in a TP program on the Fanuc controller according to the value stored in PR[1].
@@ III-B TP program
The combination of the TP program and the ROS2 hardware interface works as follows:
@@
    If R[1]==1, the TP program continuously commands the robot to move to the target joint position available in PR[1]. Using a CNT=100% motion instruction enables smooth trajectory execution.
This register-based if-then-else structure of the TP program can be extended to use other registers to command other non-motion operations. For example, one may use R[2] to command the opening/closing of a gripper through a digital input signal. Several services, such as the gripper ones in the fanuc_srvs, work according to this principle.
@@ IV Performance evaluations
To check the performance and usability of the developed hardware interface, we conducted experiments for four relevant tasks in industrial robotics: reaching a target point, tracking a trajectory, executing a collision-free trajectory, and dynamic velocity scaling. Experiments are performed with a Fanuc CRX-25iA communicating via Ethernet with a remote PC at a frequency of $25Hz$. Position and velocity measurements were recorded at the ROS2 controller’s level, which means they include communication latency and ROS2’s possible computational overheads. Comparable results were observed with a CRX-10iA_L and CRX-20iA_L robots.
@@ IV-B Trajectory tracking
To measure the tracking capabilities of the hardware interface, we streamed sinusoidal reference trajectories, $ref(t)=A\sin(2\pi ft)$, with amplitude $A=30\degree$ and frequencies $f\in\{0.1,0.25,0.5\}\,Hz$. Also, for this experiment, we requested the robot to move only the first joint. We evaluate the tracking error and the path following error.
@@
Figure 3: Results of the step responses (top), trajectory tracking (middle), and path following (bottom).
@@ IV-B2 Path following error
In this case, for the low-frequency cases ($f=\{0.1,0.25\}Hz$), the errors dramatically decrease, with averages below $1\degree$. As for the higher frequency, $f=0.5Hz$, the velocity of the joint saturates, and the resulting error remains large, leading to poor path following. The maximum error is measured when the command changes from zero to $ref(t)$. In this case, the robot should have a non-zero velocity and must adhere to physical constraints. The path errors are in the bottom rows of Table II and in Figure 3g, 3h, and 3i.
@@ IV-C Obstacle avoidance
We assessed the correct execution of collision-free trajectories via Moveit2 [10] through the experiment shown in Figure 1. The robot has to move from an initial to a goal configuration by avoiding a box-shaped obstacle. The trajectory is computed with the BiTRRT motion planner, with a velocity scaling of 0.1. The executed trajectory is in Figure 4. The plots confirm the results obtained in the previous experiments in terms of trajectory tracking and path following.
@@ IV-D Dynamic velocity scaling
[…]
We used this feature for safe human-robot interaction and human-robot collaborative transport applications. Detailed experiments of human-robot collaborative transport can be found in [14], where the proposed ros2_fanuc_interface is used in combination with the scaled_fjt_controller to adapt trajectory execution based on interaction forces.
Figure 5: Command and feedback signals with dynamic velocity scaling during trajectory execution. The speed override is set dynamically via ros topic.
@@ IV-E Limitations
The ros2_fanuc_interface does not enforce joint velocity limits. This means that large velocity commands may be clamped by the Fanuc controller. For this reason, careful attention must be paid to the robot controller’s settings, as the Fanuc controller may enforce Cartesian speed limits of around 250 mm/s in the collaborative mode. This may lead to unexpected joint speed clamping during execution.